write the first 5 terms in this sequence f(n)=5(1/2)^n

Answer by jim_thompson5910(35256)

f(n)=5(1/2)^n

f(0)=5(1/2)^(0) ... plug in n = 0

f(0)=5(0.5)^(0)

f(0)=5(1)

f(0)=5


-----------------------

f(n)=5(1/2)^n

f(1)=5(1/2)^(1) ... plug in n = 1

f(1)=5(0.5)^(1)

f(1)=5(0.5)

f(1)=2.5


-----------------------

f(n)=5(1/2)^n

f(2)=5(1/2)^(2) ... plug in n = 2

f(2)=5(0.5)^(2)

f(2)=5(0.25)

f(2)=1.25


-----------------------

f(n)=5(1/2)^n

f(3)=5(1/2)^(3) ... plug in n = 3

f(3)=5(0.5)^(3)

f(3)=5(0.125)

f(3)=0.625


-----------------------

f(n)=5(1/2)^n

f(4)=5(1/2)^(4) ... plug in n = 4

f(4)=5(0.5)^(4)

f(4)=5(0.0625)

f(4)=0.3125


-----------------------



The first five terms are: 5, 2.5, 1.25, 0.625, 0.3125